OpenStudy (anonymous):

Can someone solve 6 1/9 – 3 1/3. please show me how you got your answer.

OpenStudy (cgreenwade2000):

\[6 \frac{ 1 }{ 9 } -3 \frac{ 1 }{ 3 }\] I would start here by turning these into improper fractions. Try that out first.

OpenStudy (anonymous):

55/9 - 10/3 55/9 - 30/9 = 25/9 = 2 7/9